# Detecting Out-of-distribution through the Lens of Neural Collapse

### v5

@@ Appendix D Evaluation on DenseNet
Appendix E The Prevalence of Neural Collapse across Canonical Classification Tasks
The phenomenon of Neural Collapse, as established in the seminal work by Papyan et al. (2020) and corroborated by subsequent studies (Han et al., 2021; Mixon et al., 2020; Zhou et al., 2022; Zhu et al., 2021), widely exists across canonical classification datasets and model architectures. The prevalent occurrence of Neural Collapse forms a robust foundation for the design of our versatile OOD detectors. To this end, we review the empirical evidence of Neural Collapse across different datasets and model architectures in Figure 3, Figure 4, Figure 5, Figure 6, and Figure 7. Comparing CIFAR-10 and ImageNet behaviors with ResNet backbone in Figure 7, we note that the clustering of CIFAR-10 is more prominent than Imagenet, as indicated by a higher ratio of between-class variance to within-class covariance. Note that the figures and captions are sourced from Papyan et al. (2020). The definition and notation follow Section 3.
@@ Appendix E The Prevalence of Neural Collapse across Canonical Classification Tasks
Figure 4: (ref. Figure 3 in Papyan et al. (2020)) Classifiers and train class means approach equiangularity. In each array cell, the vertical axis shows the SD of the cosines between pairs of centered class means and classifiers across all distinct pairs of classes $c$ and $c^{\prime}$. Mathematically, denote $\cos_{\mu}(c,c^{\prime})=<\bm{\mu}_{c}-\bm{\mu}_{G},\bm{\mu}_{c}^{\prime}-\bm{\mu}_{G}>/\|\bm{\mu}_{c}-\bm{\mu}_{G}\|_{2}\|\bm{\mu}_{c}^{\prime}-\bm{\mu}_{G}\|_{2}$ and $\cos_{w}(c,c^{\prime})=<\bm{w}_{c},\bm{w}_{c}^{\prime}>/\|\bm{w}_{c}\|_{2}\|\bm{w}_{c}^{\prime}\|_{2}$, where $\{\bm{w}_{c}\}_{c=1}^{C},\{\bm{\mu}_{c}\}_{c=1}^{C}$, and $\bm{\mu}_{G}$ are as in Figure 3. We measure $\text{Std}_{c,c^{\prime}}(\cos_{\mu}(c,c^{\prime}))$ (orange) and $\text{Std}_{c,c^{\prime}}(\cos_{w}(c,c^{\prime}))$. As training progresses, the SDs of the cosines approach zero, indicating equiangularity.
@@
Figure 5: (ref. Figure 4 in Papyan et al. (2020)) Classifiers and train class means approach maximal-angle equiangularity. We plot in the vertical axis of each cell the quantities $\text{Avg}_{c,c^{\prime}}|\cos_{\mu}(c,c^{\prime})+1/(C-1)|$ (blue) and $\text{Avg}_{c,c^{\prime}}|\cos_{w}(c,c^{\prime})+1/(C-1)|$ (orange), where $\cos_{\mu}(c,c^{\prime})$ and $\cos_{w}(c,c^{\prime})$ are as in Figure 4. As training progresses, the convergence of these values to zero implies that all cosines converge to $-1/(C-1)$. This corresponds to the maximum separation possible for globally centered, equiangular vectors.
@@
Figure 6: (ref. Figure 5 in Papyan et al. (2020)) Classifier converges to train class means. The formatting and technical details are as described in Section 3. In the vertical axis of each cell, we measure the distance between the classifiers and the centered class means, both rescaled to unit norm. Mathematically, denote $\tilde{\bm{M}}=\dot{\bm{M}}/\|\dot{\bm{M}}\|_{F}$ where $\dot{\bm{M}}=[\bm{\mu}_{c}-\bm{\mu}_{G},c=1,....,C]\in\mathtt{R}^{P\times C}$ is the matrix whose columns consist of the centered train class means; denote $\tilde{\bm{W}}=\bm{W}/\|\bm{W}\|_{F}$ where $\bm{W}\in\mathtt{R}^{C\times P}$ is the last-layer classifier of the network. We plot the quantity $\|\tilde{\bm{W}}^{T}-\tilde{\bm{M}}\|_{F}^{2}$ on the vertical axis. This value decreases as a function of training, indicating that the network classifier and the centered-means matrices become proportional to each other (self-duality).
@@
Figure 7: (ref. Figure 6 in Papyan et al. (2020)) Training within-class variation collapses. In each array cell, the vertical axis (log scaled) shows the magnitude of the between-class covariance compared with the within-class covariance of the train activations. Mathematically, this is represented by $\text{Tr}(\Sigma_{W}\Sigma_{B}^{+}/C)$ where $\text{Tr}(\cdot)$ s the trace operator, $\Sigma_{W}$ is the within-class covariance of the last-layer activations of the training data, $\Sigma_{B}$ is the corresponding between-class covariance, $C$ is the total number of classes, and $[\cdot]^{+}$ is Moore–Penrose pseudoinverse. This value decreases as a function of training—indicating collapse of within-class variation.

### v7

@@ 12 Evaluation on DenseNet
13 The Prevalence of Neural Collapse across Canonical Classification Tasks
The phenomenon of Neural Collapse, as established in the seminal work by Papyan et al. [32] and corroborated by subsequent studies [10, 30, 51, 53], widely exists across canonical classification datasets and model architectures. The prevalent occurrence of Neural Collapse forms a robust foundation for the design of our versatile OOD detectors. To this end, we review the empirical evidence of Neural Collapse across different datasets and model architectures in Figure 3, Figure 4, Figure 5, Figure 6, and Figure 7. Comparing CIFAR-10 and ImageNet behaviors with ResNet backbone in Figure 7, we note that the clustering of CIFAR-10 is more prominent than Imagenet, as indicated by a higher ratio of between-class variance to within-class covariance. Note that the figures and captions are sourced from [32]. The definition and notation follow Section 3.
@@ 13 The Prevalence of Neural Collapse across Canonical Classification Tasks
Figure 4: (ref. Figure 3 in [32]) Classifiers and train class means approach equiangularity. In each array cell, the vertical axis shows the SD of the cosines between pairs of centered class means and classifiers across all distinct pairs of classes $c$ and $c^{\prime}$. Mathematically, denote $\cos_{\mu}(c,c^{\prime})=<\bm{\mu}_{c}-\bm{\mu}_{G},\bm{\mu}_{c}^{\prime}-\bm{\mu}_{G}>/\|\bm{\mu}_{c}-\bm{\mu}_{G}\|_{2}\|\bm{\mu}_{c}^{\prime}-\bm{\mu}_{G}\|_{2}$ and $\cos_{w}(c,c^{\prime})=<\bm{w}_{c},\bm{w}_{c}^{\prime}>/\|\bm{w}_{c}\|_{2}\|\bm{w}_{c}^{\prime}\|_{2}$, where $\{\bm{w}_{c}\}_{c=1}^{C},\{\bm{\mu}_{c}\}_{c=1}^{C}$, and $\bm{\mu}_{G}$ are as in Figure 3. We measure $\text{Std}_{c,c^{\prime}}(\cos_{\mu}(c,c^{\prime}))$ (orange) and $\text{Std}_{c,c^{\prime}}(\cos_{w}(c,c^{\prime}))$. As training progresses, the SDs of the cosines approach zero, indicating equiangularity.
@@
Figure 5: (ref. Figure 4 in [32]) Classifiers and train class means approach maximal-angle equiangularity. We plot in the vertical axis of each cell the quantities $\text{Avg}_{c,c^{\prime}}|\cos_{\mu}(c,c^{\prime})+1/(C-1)|$ (blue) and $\text{Avg}_{c,c^{\prime}}|\cos_{w}(c,c^{\prime})+1/(C-1)|$ (orange), where $\cos_{\mu}(c,c^{\prime})$ and $\cos_{w}(c,c^{\prime})$ are as in Figure 4. As training progresses, the convergence of these values to zero implies that all cosines converge to $-1/(C-1)$. This corresponds to the maximum separation possible for globally centered, equiangular vectors.
@@
Figure 6: (ref. Figure 5 in [32]) Classifier converges to train class means. The formatting and technical details are as described in Section 3. In the vertical axis of each cell, we measure the distance between the classifiers and the centered class means, both rescaled to unit norm. Mathematically, denote $\tilde{\bm{M}}=\dot{\bm{M}}/\|\dot{\bm{M}}\|_{F}$ where $\dot{\bm{M}}=[\bm{\mu}_{c}-\bm{\mu}_{G},c=1,....,C]\in\mathtt{R}^{P\times C}$ is the matrix whose columns consist of the centered train class means; denote $\tilde{\bm{W}}=\bm{W}/\|\bm{W}\|_{F}$ where $\bm{W}\in\mathtt{R}^{C\times P}$ is the last-layer classifier of the network. We plot the quantity $\|\tilde{\bm{W}}^{T}-\tilde{\bm{M}}\|_{F}^{2}$ on the vertical axis. This value decreases as a function of training, indicating that the network classifier and the centered-means matrices become proportional to each other (self-duality).
@@
Figure 7: (ref. Figure 6 in [32]) Training within-class variation collapses. In each array cell, the vertical axis (log scaled) shows the magnitude of the between-class covariance compared with the within-class covariance of the train activations. Mathematically, this is represented by $\text{Tr}(\Sigma_{W}\Sigma_{B}^{+}/C)$ where $\text{Tr}(\cdot)$ s the trace operator, $\Sigma_{W}$ is the within-class covariance of the last-layer activations of the training data, $\Sigma_{B}$ is the corresponding between-class covariance, $C$ is the total number of classes, and $[\cdot]^{+}$ is Moore–Penrose pseudoinverse. This value decreases as a function of training—indicating collapse of within-class variation.